OpenStudy (anonymous):

Find the volume of a triangular pyramid with base area 21 ft2 and height 8 ft. Find the result to the nearest tenth

OpenStudy (anonymous):

V=1/3*a*h a=area of base h=height V=1/3*21*8

OpenStudy (anonymous):

oh thank you so much I was putting too much into it and making it more difficult :)